The sum of two numbers is 24. Seven times the smaller number is the same as 5 times the larger number. Find the smaller number.
Answer by rchill(405)

Let x represent the smaller number, then 7 times the smaller number is 7x. Because x added to the larger number is 24, let 24-x represent the larger number. "5 times the larger number" is now represented as 5(24-x), and now our final equation is 7x=5(24-x). That expands to 7x=120-5x. Adding 5x to both sides yields 12x=120, and dividing both sides by 12 gives us x=10. So that means the larger number must be 14 (24-10=14).